# Prompt-based mental health screening from social media text

### v1

@@ Abstract
This article presents a method for prompt-based mental health screening from a large and noisy dataset of social media text. Our method uses GPT 3.5. prompting to distinguish publications that may be more relevant to the task, and then uses a straightforward bag-of-words text classifier to predict actual user labels. Results are found to be on pair with a BERT mixture of experts classifier, and incurring only a fraction of its computational costs.
@@ 1 Introduction
Textual representation models used in NLP applications have changed dramatically in recent years, moving on from simple token counts (e.g., bag-of-words) to transformer-based and large language models (LLMs) such as GPT [^1] https://platform.openai.com/docs/models and Bloom [^2] https://huggingface.co/bigscience/bloom, among many others. In particular, LLMs have successfully circumvented the need for labelled training data entirely, in so-called prompt-based methods that are now mainstream in the field, and have been shown to obtain higher accuracy in a wide range of tasks.
@@ 2 Related work
Depression detection may be seen as an instance of author profiling [Rangel et al. 2020, Flores et al. 2022, Pavan et al. 2023] from social media text. Recent studies in the field are summarised in Table 1, categorised by text genre (Reddit, Twitter), text model (b=bag of words, BERT, e=embeddings, s=sentiment), and methods.
@@
Finally, we notice that, with the exception of the work based on the SetembroBR [dos Santos et al. 2023a, dos Santos et al. 2020] corpus to be used in the present work, all of the above studies are devoted to the English language, and none make use of prompt-based methods to query a LLM for depression directly.
@@ 3 Method
As in [Cohan et al. 2018, Parapar et al. 2022] and others, our approach to depression detection relies on a dataset of social media (in our case, Twitter/X) texts that have been published by either individuals who self-reported a depression diagnosis, or by a random (control) group. Thus, the task at hand represents a binary classification problem intended to distinguish individuals who will most likely receive a depression diagnosis (called Diagnosed class) in the future from the general population (called Control class), and not to distinguish depressed from non-depressed individuals per se.
@@ 3.2 Approach
Screening for depression from social media gives rise to the questions of how to handle a large number of publications, most of which unlikely to be relevant to the task. To this end, we envisaged a prompt-based approach called Prompt.Bow that relies on GPT 3.5 to enrich an otherwise standard text classifier.
First, GPT 3.5. prompting is used to assess every tweet in our data according to relevance for mental health prediction. The (English-translated) prompt in question, adapted from the clinical description of depression in [American Psychiatric Association 2013], is shown in Figure 1.
@@
By submitting this prompt to each of the 19.42 million tweets in the corpus, the data was categorised as having high (1), medium (2) or low (3) relevance to mental health. Table shows the label distribution across the training portion of the corpus.
Table 3: Training data distribution according to relevance for depression.
@@
More importantly, however, when prompting an LLM for mental health we are largely focusing on semantics, that is, on symptoms and other well-known clinical signs of depression, but this is not to say that other factors can or should be overlooked. For instance, our current prompt allows the model to pinpoint a wide range of publications that may suggest, e.g., eating disorders or negative language use, both of which known to be related to depression, cf. [American Psychiatric Association 2013], but it does not explicitly account more fine-grained linguistic indicators of depression such as the use of first person pronouns [Trifu et al. 2017], absolute terms [Al-Mosaiwi and Johnstone 2018] and others. These indicators, which may be present in any publication of low or high relevance to depression alike, are also important predictors of mental health statuses, and no data point should in principle be discarded solely based on the LLM output.
As a means to keep the full train data available to the classifier while distinguishing between more and less relevant messages (which clearly show different distributions across classes in Table 3), Prompt.Bow uses the category labels provided by the LLM to split the training data into high, medium, and low relevance subsets, from which we create three individual bag-of-words vectors to be concatenated as discussed below.
Each of the three vectors consists of TF-IDF counts of texts labelled as high/medium/low by the GPT prompt method, and are further reduced using univariate feature selection using F1 as a score function. The final ‘high’ and ‘medium’ vectors were reduced to $k=$ 6000 features each, and the final ‘low’ vector was reduced to $k=$ 3000 features, all of which concatenated as a single 15000-word vector.
In addition to distinguishing between high-, medium- and low-relevance messages, we use the information provided by the LLM ALSO to HELP capture message order, the underlying assumption being that certain patterns (e.g., a series of consecutive ‘highly relevant’ messages) may be indicative of depression. To this end, we created a bigram model of high/medium/low labels only, reduced to its $k=$ 40000 most relevant bigram features by performing F1 univariate feature selection, and then appended to the previous 15000-word vector.
The resulting vector is taken as the input to a standard logistic regression classifier. This choice is motivated by the observation that much of the deep (e.g., semantics-driven) language processing had already been performed by the LLM, and that a simple text classifier should suffice for user label prediction.
@@ 4 Evaluation
Table 4 shows that results remain close, with a small advantage for the Prompt.Bow approach over BERT mixture of experts. However, by relying on an input provided by the pre-trained LLM, we notice that these results were obtained by using a computationally inexpensive classifier model, which represents a stark contrasts to computationally-intensive BERT mixture of experts supervision.
@@ 5 Final remarks
This work presented an experiment in prompt-based mental health screening from a large, noisy corpus of social media publications that relies on LLM prompting to distinguish publications that may be more relevant to the task, and then uses a straightforward bag-of-words text classifier to predict actual user labels. This was shown to obtain competitive results if compared to a BERT-based classifier architecture that represents the SOTA in the present setting, but incurring only a fraction of its computational costs.

### v2

@@ Abstract
This article presents a method for prompt-based mental health screening from a large and noisy dataset of social media text. Our method uses GPT 3.5. prompting to distinguish publications that may be more relevant to the task, and then uses a straightforward bag-of-words text classifier to predict actual user labels. Results are found to be on pair with a BERT mixture of experts classifier, and incurring only a fraction of its training costs.
@@ 1 Introduction
Textual representation models used in NLP applications have changed dramatically in recent years, moving on from simple token counts (e.g., bag-of-words) [Pereira and Paraboni 2007] to transformer-based and large language models (LLMs) such as GPT [^1] https://platform.openai.com/docs/models and Bloom [^2] https://huggingface.co/bigscience/bloom, among many others. In particular, LLMs have successfully circumvented the need for labelled training data entirely, in so-called prompt-based methods that are now mainstream in the field, and have been shown to obtain higher accuracy in a wide range of tasks.
@@ 2 Related work
Depression detection may be seen as an instance of author profiling [Rangel et al. 2020, Flores et al. 2022, Pavan et al. 2023, Pavan et al. 2020] from social media text. Recent studies in the field are summarised in Table 1, categorised by text genre (Reddit, Twitter), text model (b=bag of words, BERT, e=embeddings, s=sentiment), and methods.
@@
Finally, we notice that, with the exception of the work based on the SetembroBR [dos Santos et al. 2023a, dos Santos et al. 2020] corpus to be used in the present work, all of the above studies are devoted to the English language and, to the best of our knowledge, none make use of prompt-based methods to query a LLM for depression directly.
@@ 3 Method
As in [Cohan et al. 2018, Parapar et al. 2022] and others, our approach to depression detection relies on a dataset of social media (in our case, Twitter/X) texts that have been published by either individuals who self-reported a depression diagnosis, or by a random (control) group. Thus, the task at hand represents a binary classification problem intended to distinguish individuals who will most likely receive a depression diagnosis (called Diagnosed class) in the future from the general population (called Control class), and not to distinguish depressed from non-depressed individuals per se [^3] In fact, the data does not convey any guaranteed ‘non-depressed’ individuals [dos Santos et al. 2023a]..
@@ 3.1 Data
In the present work, we follow the standard train/test split provided by the corpus, as described in [dos Santos et al. 2023a].
@@ 3.2 Approach
Screening for depression from social media gives rise to the questions of how to handle a large number of publications, most of which unlikely to be relevant to the task. To this end, we envisaged a prompt-based approach called Prompt.Bow that relies on GPT 3.5 to enrich an otherwise standard text classifier. These two steps - prompting and classification - are described individually as follows.
First as a means to identify messages that are potentially related to mental health, we use GPT 3.5. prompting to assess a random sample of 30,000 tweets. The (English-translated) prompt in question, adapted from the clinical description of depression in [American Psychiatric Association 2013], is shown in Figure 1.
@@
By submitting this prompt to each of the 30,000 sample tweets, the data was categorised as having high (1), medium (2) or low (3) relevance to mental health. These GPT-labelled data was then taken as an input to train a T5 classifier [Raffel et al. 2020] to label the entire 19.42-million tweets in the corpus. Table shows the label distribution across the training portion of the data.
Table 3: Training data distribution according to relevance for depression. Tweets and tokens are shown in thousand units.
@@
More importantly, however, when prompting an LLM for mental health we are largely focusing on semantics, that is, on symptoms and other well-known clinical signs of depression. For instance, our current prompt allows the model to pinpoint a wide range of publications that may suggest, e.g., eating disorders or negative language use, both of which known to be related to depression, cf. [American Psychiatric Association 2013], but this is not to say that other factors can or should be overlooked.
In particular, we notice that LLM prompting may not explicitly account for more fine-grained linguistic indicators of depression such as the use of first person pronouns [Trifu et al. 2017], absolute terms [Al-Mosaiwi and Johnstone 2018] and other lexical factors (e.g., denoting emotion as in hate speech, cf. [da Silva et al. 2020]). These indicators, which may be present in any publication of low or high relevance to depression alike, are also important predictors of mental health statuses, and no data point should in principle be discarded solely based on the LLM output.
As a means to keep the full train data available to the classifier whilst distinguishing between more and less relevant messages (which clearly show different distributions across classes in Table 3), Prompt.Bow uses the category labels provided by the LLM to split the training data into high, medium, and low relevance subsets, from which we create three individual bag-of-words vectors (to be concatenated as discussed below). In other words, the training texts are split into three (low/medium/high relevance) categories according to the previous GPT prompt method, and we build an individual BoW model from each of these three subsets.
The resulting vectors are further reduced using univariate feature selection using F1 as a score function. The final ‘high’ and ‘medium’ vectors were reduced to $k=$ 6,000 features each, and the final ‘low’ vector was reduced to $k=$ 3,000 features, all of which concatenated as a single 15,000-word vector.
In addition to distinguishing between high-, medium- and low-relevance messages in this way, we use the information provided by the LLM also to help capture message order, the underlying assumption being that certain patterns (e.g., a series of consecutive ‘highly relevant’ messages) may be indicative of depression. To this end, we created a bigram model of high/medium/low labels only, that is, representing sequences of ‘high’, ‘medium’ and ‘low’ labels only (and not text). This was further reduced to its $k=$ 40,000 most relevant (bigram) features by performing F1 univariate feature selection, and then appended to the previous 15,000-word vector.
The resulting vector - a combination of three text models with different degrees of relevance to mental health, and a model that captures sequences of relevant messages - is taken as the input to a standard logistic regression classifier. This choice is motivated by the observation that much of the deep (e.g., semantics-driven) language processing had already been performed by the LLM, and that a simple text classifier should suffice for user label prediction.
@@ 4 Evaluation
Table 4 shows that results remain close, with a small advantage for the Prompt.Bow approach over BERT mixture of experts. However, by relying on an input provided by the pre-trained LLM, we notice that these results were obtained by using a computationally inexpensive classifier model (that is, leaving aside the costs of pre-training the LLM in the first place), which represents a stark contrasts to computationally-intensive BERT mixture of experts supervision.
@@ 5 Final remarks
This work presented an experiment in prompt-based mental health screening from a large, noisy corpus of social media publications that relies on LLM prompting to distinguish publications that may be more relevant to the task, and then uses a straightforward bag-of-words text classifier to predict actual user labels. This was shown to obtain competitive results if compared to a BERT-based classifier architecture that represents the SOTA in the present setting, but incurring only a fraction of its training costs.